OpenStudy (anonymous):

1/2= 14/(6x-2)-6/(3x-1)

OpenStudy (anonymous):

http://myalgebra.com/algebra_solver.aspx Use this link and plug in your problem. its a lifesaver

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

It's important to know why it works. Here you can see that 6x-2 is a multiple of 3x-1

hero (hero):

And just to add to that, we could also have re-written the fraction like so \(\large\frac{1}{2} = \frac{2(7)}{2(3x-1)} - \frac{6}{3x-1}\) From there we would see that the second fraction can be simplified leaving \(\large\frac{1}{2} = \frac{7}{3x-1} - \frac{6}{3x-1}\) And combing the fractions on the right side would yield \(\large\frac{1}{2} = \frac{7 - 6}{3x-1} \) Which further simplifies to \(\large\frac{1}{2} = \frac{1}{3x-1} \) Which since the numerators are the same, the denominators would also have to be the same in order for the two fractions to be equal thus: 2 = 3x - 1 If we solve for x we get: 2 + 1 = 3x 3 = 3x 3/3 = x 1 = x